OpenStudy (izzy18970):

What is the relationship between salt concentration and density? Please help

OpenStudy (aaronq):

The definition of density is \(\sf density=\dfrac{mass}{volume}\) so if you add 10 g of salt to 1 L of water you have \(d=\dfrac{10~g}{1000~mL}=0.01~g/mL\) so if you add 100 g of salt to 1 L of water you have \(d=\dfrac{100~g}{1000~mL}=0.1~g/mL\) so if you add 100 g of salt to 1 L of water you have \(d=\dfrac{1000~g}{1000~mL}=1~g/mL\) Do you see the trend?

OpenStudy (izzy18970):

I understand it now thank you so much.

OpenStudy (aaronq):

no problem, glad i could help